# Revisiting Scattering Enhancement from the Aharonov-Bohm Effect

### v1

@@ I Introduction
The interaction of dynamical matter with defects is important in the study of QFT given the varied roles defects can play. For example, defect operators can serve as order parameters such as how the expectation value of the ‘t Hooft operator can detect confinement [1] or have big phenomenological implications such as through the Callan-Rubakov effect in the interaction of monopole lines with matter [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
@@ II Discrete Gauge Theory and the Aharonov-Bohm Effect
Now consider the fermion field in the presence of an ${\cal S}_{q}$ operator wrapping the $z$-axis on ${\mathbb{R}}^{3}\times{\mathbb{R}}_{t}$. Due to the equation of motion
$\begin{split}\frac{dA_{1}}{2\pi}=\frac{q}{N}\delta^{(2)}(x)\end{split}$ (8)
$A_{1}$ must have holonomy $e^{i\oint_{\gamma}A_{1}}=e^{\frac{2\pi iq}{N}}$ for all loops $\gamma$ that wind once around the origin in the $x-y$ plane. Since the fermion $\Psi(x)$ has charge 1 under the $A_{1}$ gauge group, the field takes value in sections of the spinor bundle tensored with a line bundle with connection $A_{1}$ over a spacetime where we have excised an infinitesimal tubular region around the string operator. Since the holonomy of the line bundle is fixed by $A_{1}$, the fermion field $\Psi(x)$ must have the periodicity condition
@@
where $\theta$ is the polar angle in the $x-y$ plane. This is the Aharonov-Bohm effect.
Another way to see that the operator ${\cal S}_{q}$ both turns on the background gauge field and causes $\Psi(x)$ to be multi-valued is to use the language of generalized symmetries. By analyzing the Lagrangian in (6), we see that the gauge field is shifted $A\mapsto A+\alpha\,\delta(\sigma)$ by the operator [^1] Note that here we are working with singular transformations which are a typical feature of the operator formalism.
@@
where $\partial\sigma=\Sigma$. Additionally, the operator $e^{i\alpha\int_{\sigma}\ast j_{\Psi}}$ induces a phase jump $e^{i\alpha}$ of the $\Psi$ field across the manifold $\sigma$. Since ${\cal S}_{q}(\Sigma)=V_{\alpha}(\Sigma;\sigma)\,e^{-\frac{2\pi iq}{N}\int_{\sigma}\ast j_{\Psi}}$, it is clear that ${\cal S}_{q}$ both shifts the gauge field and causes $\Psi(x)$ to be multi-valued. Said differently, the insertion of the operator ${\cal S}_{q}$ in the path integral can be absorbed into the action only by both shifting the gauge field and by performing a field redefinition that makes $\Psi(x)$ multi-valued.
@@ Acknowledgements
The authors would like to thank Ken Intriligator, Aneesh Manohar, Oren Bergman, Ben Grinstein, Sungwoo Hong, John March-Russell, LianTao Wang, and Seth Koren for helpful discussions and related collaborations. TDB is supported by Simons Foundation award 568420 (Simons Investigator) and award 888994 (The Simons Collaboration on Global Categorical Symmetries). JSG acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [PGS D - 587640] and U.S. Department of Energy (DOE) award number DE-SC0009919.

### v2

@@ I Introduction
The interaction of dynamical matter with defects is important in the study of QFT given the varied roles defects can play. For example, defect operators can serve as order parameters such as how the expectation value of the ‘t Hooft operator can detect confinement [1] or have big phenomenological implications such as through the Callan-Rubakov effect in the interaction of monopole lines with matter [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13].
@@ II Discrete Gauge Theory and the Aharonov-Bohm Effect
Now consider the fermion field in the presence of an ${\cal S}_{q}(\Sigma)$ operator wrapping the $z$-axis on ${\mathbb{R}}^{3}\times{\mathbb{R}}_{t}$. In the partition function, $A_{1}$ is a $U(1)$ gauge field with $\oint\frac{dA_{1}}{2\pi}\in{\mathbb{Z}}$ and $\Psi$ is a single-valued field. After inserting ${\cal S}_{q}$ into the path integral, the $B_{2}$ field can be integrated out by shifting $A_{1}\mapsto A_{1}^{\prime}=A_{1}+\frac{2\pi q}{N}\frac{d\theta}{2\pi}$ for $\partial\sigma=\Sigma$ and modifying the periodicity condition for $\Psi$ so that it obeys
@@
where $\theta$ is the polar angle in the $x-y$ plane. This eliminates ${\cal S}_{q}$ and then integrating out $B_{2}$ restricts $A_{1}^{\prime}$ to be a ${\mathbb{Z}}_{N}$ gauge field while at the same time inducing an Aharonov-Bohm phase on $\Psi$.
Another way to see that the operator ${\cal S}_{q}$ both turns on the gauge field and causes $\Psi(x)$ to be multi-valued is to use the operator formalism of generalized symmetries. Analyzing the Lagrangian in (6) shows that the gauge field is shifted $A\mapsto A+\alpha\,\delta(\sigma)$ by the operator [^1] Note that here we are working with singular transformations which are a typical feature of the operator formalism.
@@
where $\partial\sigma=\Sigma$ and the operator $e^{i\alpha\int_{\sigma}\ast j_{\Psi}}$ induces a phase jump $e^{i\alpha}$ of the $\Psi$ field across the manifold $\sigma$. Since ${\cal S}_{q}(\Sigma)=V_{\frac{2\pi q}{N}}(\Sigma;\sigma)\,e^{-\frac{2\pi iq}{N}\int_{\sigma}\ast j_{\Psi}}$, it is clear that ${\cal S}_{q}$ both shifts the gauge field and causes $\Psi(x)$ to be multi-valued.
@@ Acknowledgements
The authors would like to thank Ken Intriligator, Aneesh Manohar, Oren Bergman, Ben Grinstein, Sungwoo Hong, John March-Russell, LianTao Wang, Seth Koren, and Prateek Agrawal for helpful discussions and related collaborations. TDB is supported by Simons Foundation award 568420 (Simons Investigator) and award 888994 (The Simons Collaboration on Global Categorical Symmetries). JSG acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) [PGS D - 587640] and U.S. Department of Energy (DOE) award number DE-SC0009919. JSG also thanks the Galileo Galilei Institute for Theoretical Physics for their hospitality during the completion of this work.